### v1

# First Demonstration of Underground Muon Imaging at an Archaeological Site in Ancient Jerusalem

Abstract
We present a novel underground imaging system that utilizes cosmic-ray muons to explore the subsurface environment at the City of David archaeological site in ancient Jerusalem. This report details the initial findings from measurements conducted at a large cistern, commonly called “Jeremiah’s cistern” (referenced in Jeremiah 38:6). The system aims to locate and map hidden voids and structural anomalies within the overburden. Our primary outcome is the derivation of the angular ground depth, which serves as a proxy for understanding the integrated density distribution of the overburden. This work represents a significant interdisciplinary effort to deepen our understanding of this historically important site.
I Introduction
Muon tomography is a non-invasive imaging technique that exploits cosmic-ray muons to probe the density distribution within a target material Alvarez et al. (1970); Nagamine (2008); Schouten (2018). As muons traverse matter, density variations perturb their trajectories, enabling the reconstruction of subsurface structures by analyzing their altered paths. This technique has recently been applied in geological and archaeological investigations Tanaka et al. (2003); Schouten (2018); D’Errico et al. (2020).
The south-eastern ridge (a.k.a. the City of David) in ancient Jerusalem, with stratified deposits dating back to the second and first millennia BCE Reich (2011), provides an excellent test bed for muon tomography. This work describes the first operational test of our custom-developed muon detector deployed at Jeremiah’s cistern, a site adjacent to ongoing excavations. Our objectives include (i) demonstrating the efficacy of muon tomography for archaeological imaging, (ii) performing flux measurements, (iii) deriving integrated opacity and ground depth, and (iv) preparing for a second campaign, where the detector is placed underneath the City of David near the Gihon spring.
The cistern is a large rock-hewn installation. A picture of the cistern is shown in Fig. 1. The feature, approximately six meters deep with a narrow shaft and a bell-shaped lower chamber, exhibits typical water system characteristics and was likely used for water storage or containment. The cistern date is yet unclear. Nevertheless, it is located at the top of the ridge, adjacent to the Large Stone Structure and the rock cut moat, two major structures interpreted by some as a royal construction project, indicating that the installation may have served as a prominent building complex during the First Temple period Mazar (2007, 2020); Gadot et al. (2023).
Figure 1: Photograph of Jeremiah’s cistern, the bell-shaped water facility underneath the City of David visitor center.
II Experimental Setup and Methods
II.1 Detector Design and Calibration
The muon detector is a compact, scintillator-based telescope optimized for underground applications. It comprises four detection layers containing 23 extruded plastic scintillator bars, 40 cm long with a triangular cross-section of $3.3\times 1.7\,\text{cm}^{2}$. The scintillators are fabricated from polystyrene doped with PPO and POPOP and coated with titanium oxide (see Ref. Pla-Dalmau, Bross, and Mellott (2001)). When traversed by a muon, scintillation light is produced and captured by wavelength-shifting (WLS) fibers (Saint-Gobain BCF-91A). These fibers shift the blue light to green and guide it to Silicon Photomultipliers (SiPMs)Onsemi mounted on PCBs. Each fiber extends an additional 50 cm beyond the scintillator to mitigate attenuation losses. A detailed description of the detector is given in Ref. Benhammou et al. (2022).
Calibration of the detector involved measuring individual scintillator responses, gain matching for the SiPMs, and determining the effective detection surface $S_{\mathrm{eff}}(\tau_{x},\tau_{y})$. Analysis of initial data revealed that trigger thresholds for muon hit identification were initially set too high, resulting in a detection efficiency of approximately 30%. These thresholds have since been optimized and are now correctly configured for current data acquisition.
II.2 Data Acquisition System
The detector employs the CAEN DT5550W data acquisition system based on WeeROC ASICs CAEN . A mezzanine card with four 32-channel Citiroc ASICs, combined with an FPGA (Xilinx XC7K160T Xilinix ), processes the signals. A 14-bit, 80 MS/s ADC digitizes the signals in real time. The system is powered by 128 bias voltage lines (20–35 V) and housed in a dedicated cooling box to ensure thermal stability.
Figure 2 shows a photograph of the detector during the installation and the commissioning in Jeremiah’s cistern.
Figure 2: The muon detector during its commissioning inside Jeremiah’s cistern.
II.3 Muon Trajectory Determination
Each cosmic-ray muon is expected to leave a trace in all four scintillator layers. For sub-pixel hit resolution, the photon yield in a scintillator is proportional to the muon’s path length within the bar. When a muon traverses the interface between two adjacent triangular scintillator bars, it produces signals $n$ and $N$ in the two bars, respectively. Figure 3 illustrates the geometry of this process (taken from the description of the detector given in Ref. Benhammou et al. (2022)).
Figure 3: Illustration of a muon passing through two adjacent triangular scintillator bars (Ref. Benhammou et al. (2022)). The signals $n$ and $N$ are proportional to the muon path lengths in the respective bars, enabling a refined determination of the hit position $X$ from the bar vertex (with side length $a$).
Assuming the muon crosses the interface at a distance $X$ from the vertex of a bar with side length $a$, the ratio of the signals satisfies:
$\frac{n}{N}=\frac{X}{a-X}.$ (1)
This relation lets us estimate $X$ with subcentimeter precision. Since the signals $n$ and $N$ follow Poisson statistics ($\Delta n=\sqrt{n}$ and $\Delta N=\sqrt{N}$), the uncertainty in $X$ is computed as:
$\delta X^{2}=\left(\frac{\partial X}{\partial N}\Delta N\right)^{2}+\left(\frac{\partial X}{\partial n}\Delta n\right)^{2},$ (2)
which yields:
$\delta X=\frac{a}{N+n}\sqrt{\frac{n\cdot N}{N+n}}.$ (3)
This sub-pixel resolution is essential for achieving the overall angular resolution of approximately 20 mrad.
For overall trajectory reconstruction, the refined hit positions from each layer are correlated. Denote the hit positions in the top and bottom layers as $x_{\text{top}}$ and $x_{\text{bottom}}$ (and similarly for $y$). The differences $\Delta x=x_{\text{bottom}}-x_{\text{top}}$ and $\Delta y=y_{\text{bottom}}-y_{\text{top}}$ are computed over a known vertical separation $\Delta z$. The muon incident angles are then given by:
$\tau_{x}=\tan(\theta_{x})=\frac{\Delta x}{\Delta z},\quad\tau_{y}=\tan(\theta_{y})=\frac{\Delta y}{\Delta z}.$ (4)
II.4 Flux Measurement and Angular Ground Depth Estimation
Muon trajectories are binned in the angular space $(\tau_{x},\tau_{y})$ to form the count matrix $C_{\mu}(\tau_{x},\tau_{y})$, and Poisson statistics govern the uncertainty in each bin.
The muon rate per angular bin is defined as:
$R_{\mu}(\tau_{x},\tau_{y})=\frac{C_{\mu}(\tau_{x},\tau_{y})}{T_{\mathrm{det}}},$ (5)
with $T_{\mathrm{det}}$ being the exposure time. The differential muon flux is then calculated by:
$F_{\mu}(\tau_{x},\tau_{y})=\frac{R_{\mu}(\tau_{x},\tau_{y})}{S_{\mathrm{eff}}(\tau_{x},\tau_{y})\,J},$ (6)
where $S_{\mathrm{eff}}(\tau_{x},\tau_{y})$ is the effective detection surface (which accounts for the geometric acceptance) and $J=\cos^{3}\theta$ is the Jacobian for the transformation from $(\tau_{x},\tau_{y})$ to spherical angles $(\theta,\phi)$.
II.5 Estimation of Integrated Opacity
Lookup tables derived from muon transport simulations are used to invert the measured flux and estimate the angular ground depth $D(\tau_{x},\tau_{y})$, which reflects the effective overburden thickness. This result is presented as the primary outcome of our study.
The muon flux is reduced while traversing through the ground because the muons lose energy while penetrating until they stop. The muon flux is related to the integrated density (or opacity) $\rho L$ (in g cm-2) along the muon path via energy-dependent attenuation:
$\frac{dE}{dX}=-a(E)-b(E)\cdot E$ (7)
Here, $a(E)$ represents nearly constant ionization losses (approximately $2\,\mathrm{MeV\,cm^{2}/g}$, per the Bethe-Bloch formula) and $b(E)$ represents radiative losses (including bremsstrahlung, pair production, and photonuclear interactions) that become significant at higher energies Groom, Mokhov, and Striganov (2001); Navas et al. (2024). The integrated opacity is obtained by solving:
$\int_{E_{\mathrm{min}}}^{E_{\mathrm{GL}}}\frac{dE}{a(E)+b(E)\,E}=\rho L,$ (8)
where $E_{\mathrm{GL}}$ is the muon energy at ground level and $E_{\mathrm{min}}$ is the threshold energy for muon penetration. Detailed muon transport simulations yield lookup tables correlating the measured flux with $\rho L$.
For this, we simulated the expected muonic flux using the parameterization for the vertical intensity shown in  Reyna (2006) and  Bugaev et al. (1998). Specifically, the muon intensity at any zenith angle is expressed as
$I(p_{\mu},\theta)=\cos^{3}\theta\,I_{v}\bigl{(}p_{\mu}\cos\theta\bigr{)},$ (9)
where the vertical muon intensity $I_{v}(p_{\mu})$ is given by
$I_{v}(p_{\mu})=c_{1}\,p_{\mu}^{-1{\bigl{(}c_{2}+c_{3}\log_{10}(p_{\mu})+c_{4}\log_{10}^{2}(p_{\mu})+c_{5}\log_{10}^{3}(p_{\mu})\bigr{)}}}$ (10)
Using the best-fit coefficients $c_{1}=0.00253$, $c_{2}=0.2455$, $c_{3}=1.288$, $c_{4}=-0.2555$, and $c_{5}=0.0209$, this combined parameterization accurately reproduces the observed muon intensities over a broad momentum range. Equations (9) and (10) were incorporated into our simulation framework to predict the angular and momentum distribution of muons at the surface, which was subsequently used to estimate the expected muonic flux for every angle and depth.
II.6 Uncertainty and Error Analysis
Accurate error quantification is critical because it is essential for accurately reconstructing deviations in the ground. Reliable uncertainty estimation directly affects the validity of the results. The main sources of uncertainty include:
  • Detector Efficiency: The overall efficiency (less than 20%)—due to high trigger thresholds and non-optimal trigger logic—introduces systematic bias.
  • Atmospheric Modeling: Our simulation used sea-level muon flux models; however, Jerusalem’s altitude (approximately 650 m) yields a higher flux and slightly altered angular distribution.
  • Simulations: Limitations in the muon transport simulations, mainly due to insufficient modeling of the surface and terrain, contribute additional uncertainties.
  • Statistical Fluctuations: Poisson statistics dictate the uncertainty in each angular bin.
The follow-up work aims to optimize trigger thresholds, incorporate altitude corrections, and refine simulation models, which is expected to reduce major contributions to uncertainty.
III Results
After thorough calibration and extensive laboratory testing, the detector was deployed at Jeremiah’s cistern in the City of David archaeological site. Laser imaging, Detection, and Ranging (LIDAR) mapping data were integrated into our simulation framework to predict the expected muon flux under homogeneous overburden conditions.
III.1 Integrated muon rate and overall efficiency
The detector was deployed and operated at Jeremiah’s cistern for a total duration of ten days. The total muon rate $\bigl{(}\int_{\tau_{x}}\int_{\tau_{y}}R_{\mu}(\tau_{x},\tau_{y})\bigr{)}$ and the rate after the quality selection criteria that was acquired in the cistern during this time is shown in figure 4. In principle, we have used only those muon tracks that were analyzed with super resolution (i.e. were recorded in two adjacent bars in each of the four detector layers). By comparing these measured total rates to the expected one (2.45($\approx$ 2.45( ≈ 2.45 [Hz]), we estimate the total detector efficiency to be $\approx 50\%$ for the total rate but only $\approx 15\%-20\%$ for the actual muons used.
Figure 4: Muon rates during the measurement calculated in 20 minutes bins, (blue) the total muon rate, (red) muon rate after quality selection criteria.
III.2 LIDAR Mapping and Flux Distribution
Figure 5 shows a LIDAR image of the interior of Jeremiah’s cistern, clearly revealing the ventilation shafts extending toward the ground surface and the detector’s location. Figure 6 presents the $\tan\theta_{x}$ versus $\tan\theta_{y}$ distribution of the muon flux. The top panel shows simulation results, while the bottom displays the measured data. Both distributions exhibit a maximum flux near the zenith, with a distinct enhancement corresponding to the known ventilation shaft.
Figure 5: LIDAR image of the interior of Jeremiah’s cistern, highlighting the air shafts and the detector’s location.
Figure 6: $\tan\theta_{x}$ versus $\tan\theta_{y}$ distribution of muon flux: (top) simulation, (bottom) measured data.
III.3 Measured Angular Ground Depth
By applying the measured differential flux to lookup tables from muon transport simulations, we derive the angular ground depth $D(\tau_{x},\tau_{y})$, which acts as a proxy for the integrated density of the overburden. Figure 7 shows the resulting density map. This map, the primary result of our study, reveals variations in the overburden that may indicate subsurface anomalies such as voids or high-density inclusions.
Figure 7: The angular ground depth (meters) $D(\tau_{x},\tau_{y})$ derived from the transformation of the observed muon flux.
IV Discussion
The measurement demonstrates that muons can be used to map underground features. The discrepancies between the measured and expected muon flux and overburden can be attributed to several factors. First, the detector efficiency is low (less than 20%), primarily due to a suboptimal triggering scheme and excessively high thresholds for muon hit identification. This inefficiency results in significant undercounting of muon events and a reduced flux relative to simulation predictions. Optimization of the trigger logic and recalibration of the thresholds are necessary steps for improving the measurement’s sensitivity and accuracy.
Second, the simulation of the atmospheric muon flux assumed a sea-level model. Since Jerusalem is located approximately 650 m above sea level, the actual muon flux is higher, with a slightly altered angular distribution. This altitude effect introduces a systematic discrepancy in both the overall muon rate and the angular profile, affecting the accuracy of the derived overburden.
Third, while the interior part of the cistern was precisely scanned with LIDAR, the ground level was assumed to be flat and without substantial infrastructure. However, this approximation is inaccurate; the ground above Jeremiah’s cistern is not flat and has several $\mathcal{O}(meter)$ features, namely massive stone walls, large olive trees, and buildings surrounding the premise. These were not incorporated into the simulations.
Additional uncertainties stem from limitations in the muon transport simulations, environmental variations, and potential systematic uncertainties in detector calibration. The current effort focuses on incorporating altitude corrections into the simulation models, improving calibration procedures, and deploying multiple detectors to achieve three-dimensional subsurface imaging.
V Conclusion
This study has successfully demonstrated the utility of muon imaging as an effective non-invasive imaging technique for mapping underground features, specifically within the archaeological context of Jeremiah’s cistern in the City of David, Jerusalem. We could derive key insights into the subsurface density variation and overburden characteristics by deploying a custom muon detector and integrating advanced LIDAR mapping techniques. While discrepancies between the measured and expected muon flux were identified, primarily attributed to detector efficiency, altitude effects, and the inherent complexities of the ground structure, the findings pave the way for further refinement and optimization of the muon detection process. Follow-up work will address these uncertainties by enhancing calibration methods, incorporating altitude adjustments in simulations, and expanding the network of detectors for comprehensive three-dimensional imaging. Overall, this research highlights the potential of muon tomography in archaeological investigations and encourages its continued application in exploring buried structures and understanding ancient civilizations.
VI Data Availability
The data supporting this study’s findings are available from the corresponding author upon reasonable request.
Acknowledgements.
This work was supported in part by the Pazy and Cogito foundations. The site is located in the City of David in the Jerusalem Walls National Park. Y. Shalev and F. Vukosavović excavated on behalf of the IAA, Permit No. 8928.